40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3383     Entry time: Sat Aug 7 11:07:44 2010
Author: Koji 
Type: Configuration 
Category: PSL 
Subject: MC kept locked / PMC control gain reduced to +13dB 

Jenne asked us to keep th MC locked and let the seismometers happy through this weekend.
Note that the work at the control room and the desks are no problem as far as you are quiet.

Nancy told Jenne and me that she finished the work and reverted the WFS to the old state at 4:30AM.
She could not make the elog as it has been crashed.

MC and old MC WFS looks working as usual.

From 6:40AM to 9:40AM the oscillation of the PMC looks present.

At 10:30AM I reduced the gain of the PMC from +15dB to +13dB.

ELOG V3.1.3-